Grade 12Physical Chemistry

suppose it is required to make a buffer solution of pH=4,using acetic acid and sodium acetate is to be added to 1 litre of N÷10 acetic acid? dissociation constant of acetic acid=1.8×10?5 ?

lokesh palingi
12 Years agoGrade 12

1 Answer

Sunil Kumar FP
12 Years ago
applying thr formula
ph=pka+ log(salt/acid)
given ph= 4
pka=-log(1.8× 10^-5) =4.74
[acid]=.1
applt it and get the answer what you want
thanks and regards
sunil kr
askIItian faculty